Operation Summary for June 7, 2010 - June 9, 2010 |
Monday June 7 The day shift began with Operations monitoring the Tevatron (TeV) store 7874 (95.01E30) and the Recycler stash (272.1E10), with the Antiproton Source (Pbar) stacking (15.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7874 at 10:30 AM. Operators prepared the CDF and D0 collision halls for access while cryo system technicians worked on various TeV wet engines. Recycler personnel conducted beam studies prior to shot setup. Operations put the TeV into shot setup at 2 PM and established store 7877 at 3 PM with an average initial luminosity (AIL) of 372E30. During the evening shift, a Booster RF ion pump (BRF14) tripped off multiple times during the shift. Operators tuned the 750 KeV beam line. Tuesday June 8 The midnight shift began with Operations monitoring store 7877 (114.49E30) and the stash (258.2E10), with Pbar stacking (19.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators search and secured the Meson MTest section 2 enclosure and then established beam to the FTBF T-1006 MINERvA experiment at 4:17 AM. Operations terminated store 7877 at 4:49 AM, put the TeV into shot setup at 4:53 AM, and established store 7878 at 6:07 AM with an AIL of 353.63E30. Operations reported a Booster vacuum cascade at 5:51 AM. The vacuum was unrecoverable. Operators closed the vacuum valves to isolate the region. Operations turned off the Gradient Magnet Power Supply for the Booster vacuum repairs. Operations turned off the NuMI beam at 7:40 AM for an access. During the day shift, Booster personnel located the vacuum leak on the cavity for a Booster RF station (BRF3); they removed the cavity, installed a spool piece, and had the vacuum pumps on shortly after noon. Various personnel accessed enclosures MI-12A&B, the MI-65 Absorber Hall, and the Main Injector for short maintenance jobs. Pbar personnel conducted cooling studies. During the evening shift, the Booster vacuum recovered at 6:25 PM. Booster personnel made another access to repair a beam loss monitor in the tunnel. Booster resumed sending beam at 8:15 PM. Operations reported that a MI RF station (MIRF5) tripped off several times; operators adjusted its feed forward system. Wednesday June 9 The midnight shift began with Operations monitoring store 7878 (23.29E30) and the stash (78.1E10), with Pbar stacking (46.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, TeV mid-store studies began at midnight and concluded at 2:55 AM. Operators traced Recycler cooling issues to a Pelletron energy error. Operations established beam to T-977 at 4:50 AM. Operations terminated store 7878 at 5:05 AM. TeV personnel began CDF aperture scans at 5:40 AM. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 06/09/2010 email Fermilab |
| Security, Privacy, Legal | |